OpenStudy (anonymous):

use the method of substitution to solve the system of linear equations: 5a=7-b 5a=6-b

OpenStudy (anonymous):

bro you better medal me for this

OpenStudy (anonymous):

okay move everything over BUT B

OpenStudy (anonymous):

and then substitute

OpenStudy (anonymous):

tadaaaaa

OpenStudy (campbell_st):

there is no solution to this system of equations..

OpenStudy (anonymous):

no solution???

OpenStudy (campbell_st):

thats right both equations have 5a... so substitute the 2nd equation into the 1st and you get 6 - b = 7 - b add b to both sides 6 = 7 doesn't work

OpenStudy (anonymous):

ok, I am new to this site...karolinarivas5 said that I had better "Medal Him"...what does this mean??

OpenStudy (campbell_st):

don't worry about the medal.... its about understanding the process of solving your equations... do you understand what I've done in my solution..?

OpenStudy (anonymous):

Yes, I do see where you combined and added b to both sides...and that both equations has a 5...Thanks for your help!!